Calculate the number of grams of O2 required for the complete combustion of 1.50kg of C2H2:2C2H2+—> 4CO2+2H2O
1
Expert's answer
2020-10-12T13:36:23-0400

2C2H+ 5O2 → 4CO2 + 2H2O

m(C2H2) = 1.50 kg = 1.50*103 g.

Let's make a proportion according to the reaction equation:

1.50*103 g C2H— x g O;

2*26 g C2H2 — 5*16 g O2 .

"x = \frac{1.50*10^3*5*16}{2*26} = 2308 (g) \approx\ 2310\ g\ —\ m (O2)"